OpenStudy (anonymous):

a person with a black belt in karate has a fist that has a mass of 0.59kg. starting from rest, this fist attains a velocity of 7.1m/s in 0.11 sec. what is the magnitude of the average net force applied to the fist to acheive this level of performace.

OpenStudy (anonymous):

mass (m) = 0.59 kg initial velocity (Vo) = 0 m/s final velocity (Vf) = 7.1 m/s time (t) = 0.11 sec \[F = m*a\] \[F = m( V _{f} - V _{i} ) / t\] \[F = ( 0.59 kg )( 7.1 m/s - 0 m/s ) / 0.11 \sec\] \[F = 38.0818 N\] the magnitude of the average net force is 38.0818 Newtons.